OpenStudy (anonymous):

f(x)= 4x^2 -3 What is the domain if the range is 1<=f(x)<=13? Is the domain -2<=x<=2, 1<=x<=2, or -2<=x<=-1? And why?

OpenStudy (anonymous):

1<=4x^2-3<=13 4<=4x^2<=16 1<=x^2<=4 1<= x <=2

OpenStudy (anonymous):

and -2<=x<=-1

OpenStudy (anonymous):

In real numbers, 1<=x<=2 U -2<=x<=-1

OpenStudy (anonymous):

Makes sense